OpenStudy (anonymous):

given f(x)=3x^2+5x+2 and g(x)+x-1: a)find (f+g)(x) b)find (fg)(x) c)find (f*g)(x)

OpenStudy (anonymous):

c) is an open circle between f and g dont know what its called

OpenStudy (anonymous):

I am assuming g(x)=x-1 a)(f+g)(x) simply means f(x)+g(x) so we find just add the two functions, (3x^2+5x+2) + (x-1) which gives (3x^2+6x+1)

OpenStudy (anonymous):

b)(fg)(x) means that the two functions are multiplied, this means f(x)*g(x) this gives (3x^2+5x+2)(x-1) =(3x^3+2x^2-3x-2) c)open circle that is, (f o g)(x) means, f(g(x)). this means that wherever there is x in the function f, we replace it with the value of g(x).that means, in the function f, we replace x with x-1 so now we get, f(x-1).so this is 3(x-1)^2+5(x-1)+2 = 3(x^2 -2x +1)-5(x-1)+2 =3x^2-11x+10

OpenStudy (anonymous):

thankyou really appriciate it